# Are Colors Quanta of Light for Human Vision? A Quantum Cognition Study of Visual Perception

### v1

@@ Abstract
We study the phenomenon of categorical perception within the quantum measurement process. The mechanism underlying this phenomenon consists in dilating stimuli being perceived to belong to different categories and contracting stimuli being perceived to belong to the same category. We show that, due to the naturally different way in determining the distance between pure states compared to the distance between density states, the phenomenon of categorical perception is rooted in the structure of the quantum measurement process itself. We apply our findings to the situation of visual perception of colors and argue that it is possible to consider colors as light quanta for human visual perception in a similar way as photons are light quanta for physical measurements of light frequencies. In our approach we see perception as a complex encounter between the existing physical reality, the stimuli, and the reality expected by the perciever, resulting in the experience of the percepts. We investigate what that means for the situation of two colors, which we call Light and Dark, given our findings on categorical perception within the quantum measurement process.
Keywords:human vision, categorical perception, quantum measurement, Bloch sphere, quantisation, basic colors, qubits, quantum cognition
@@ 1 Introduction
We will argue that colors are quanta of light for human vision in a similar way that photons are quanta of light measured by a physical measuring device. To gather evidence for this argument, we will draw on various results obtained over the years within the research field of quantum cognition (Aerts & Aerts, 1995; Aerts & Gabora, 2005b; Busemeyer et al., 2006; Aerts, 2009a, b; Bruza & Gabora, 2009; Aerts & Sozzo, 2011; Aerts et al., 2012; Busemeyer & Bruza, 2012; Haven & Khrennikov, 2013; Khrennikov, 2014; Dalla Chiara et al., 2015; Pothos et al., 2015; Blutner & beim Graben, 2016; Moreira & Wichert, 2016; Yearsley, 2017; Aerts Arguëlles, 2018; Busemeyer et al., 2019; Surov et al., 2019; Aerts Arguëlles & Sozzo, 2020; Aerts & Beltran, 2022a), some of which also involve our own work (Aerts Arguëlles, 2018; Aerts Arguëlles & Sozzo, 2020). Nevertheless, we very deliberately want to make this article understandable to those who have not studied the investigations on which it relies, so that in addition to references to this research, and more technical portions of the article, we will also represent in an intuitive clear way the core aspects in which we frame our claim.
@@
Our argument rests on a fundamental property of human perception, which is not limited specifically to human vision, but is present in all forms of human perception, and which has its origin in a phenomenon called ‘categorical perception’ (Harnad, 1987; Goldstone & Hendrickson, 2010; Aerts & Aerts Arguëlles, 2022). Categorical perception ‘introduces a warping’ of the ‘stimulus’ that generated the perception in a very specific way. Namely, stimuli belonging to the same category are perceived in a way that they are more similar for the one perceiving than an independent of human bias measurement of the stimuli reveals. Additionally, for stimuli belonging to a different category, the reverse warping takes place, they are perceived in a way that they are more different for the one perceiving than an independent of human bias measurement of the stimuli reveals. There is a long history (Bruner & Postman, 1949; Liberman et al., 1957, 1967; Lane, 1965; Eimas et al., 1971; Lawrence, 1949; Berlin & Kay, 1969; Davies et al., 1998; Davidoff, 2001; Regier & Kay, 2009; Havy & Waxman, 2016; Hess et al., 2009; Disa et al., 2011; Sidman, 1994; Schusterman et al., 2000; Rosch, 1973; Collier at al., 1973) in which the phenomenon of categorical perception was identified incrementally, and in Section 2 we give a brief description of this genesis.
@@
In the reasoning we develop we also rely on a detailed study of the quantum measurement process, and more specifically on the way this was analyzed, using a specific measurement model, by our research group in the foundations of quantum mechanics at the Free University of Brussels (Aerts, 1986; Aerts & Sassoli de Bianchi, 2015, 2016, 2017). The mathematical model we developed is an extension of the Bloch model, where not only the pure states, the collapse probabilities and the decoherenced density states are represented, but also the measurements and the changes caused by them. We will see in the course of our analysis how this extended Bloch model is a fertile ground for identifying the phenomenon of categorical perception, and we will show that the quantum measurement model already intrinsically contains the structure that gives rise to the presence of the measurement bias of categorical perception. We will then gather the arguments that allow us to consider within the analysis we propose the basic colors as quanta of light for human visual perception in a similar way that photons are considered quanta of light for physical measurement apparatuses. For reasons of simplicity we introduce the mathematical elements for our analysis for the case of two colors that we call Light and Dark, mentioning that fundamentally the same analysis can be made for multiple colors with mathematical elements that then belong to a more than two dimensional complex Hilbert space. No other specific difficulties appear in this more dimensional situation that are not encountered in the two dimensional case, namely that of two colors and a two dimensional Hilbert space, which we consider in detail in the present article.
@@
In Section 2 we give a brief overview of the phenomenon of categorical perception with specific attention to the case of colors. In Section 3 we introduce in a detailed and self-prescriptive way, so that prior knowledge is not necessary, the quantum measurement model, paying attention to the extension of the traditional Bloch model that also includes a representation of the measurement and the changes caused by it. In Section 4, we analyze the way the quantum measurement changes distances between states and show that this way contains exactly the warping associated with categorical perception. In other words, the bias caused by categorical perception is intrinsically contained in the structure of the quantum measurement process. In Section 5 we gather the elements of the previous sections to give expression to our claim that we can consider colors as quanta of light for human visual perception.
@@ 2 Categorical Perception, Quantization and Colors
In this section we wish to bring out the phenomenon of categorical perception, because the nature of this phenomenon, along with the details of the measurement model of quantum mechanics, which our next Section 3 deals with, plays a fundamental role in our claim that colors are quanta of light for human vision. It is interesting to reflect for a moment on the emergence of what was later commonly called categorical perception because it illustrates well the role that certain goals and hypotheses can play.
@@
Around 1950, research on speech perception came into focus as work was being done on a speech device, the ‘pattern playback machine’, which was intended to make it possible to automatically convert texts into spoken form so that, for example, blind people could read with it. This required analyzing speech very thoroughly, and it is within this setting that Alvin Liberman identified the phenomenon that would become known as categorical perception. More specifically, he noticed that when generating a continuum of evenly distributed consonant-vowel syllables with endpoints reliably identified as ‘b’, ‘d’ and ‘g’, there is a point of rapid decrease in the probability of hearing the sound as a ‘b’ to hearing it as ‘d’. At a later point, there is a rapid switch from ‘d’ to ‘g’ (Liberman et al., 1957). Liberman formulated an original hypothesis by which he wished to explain why people perceive an abrupt change between ‘b’ and ‘p’ in the way speech sounds are heard in contrast to what happens with a synthetic morphing device that produces the sounds with a continuous transition. His hypothesis was that this phenomenon is due to a limitation of the human speech apparatus which, due to the muscular nature of its construction, would be unable to produce continuous transitions. Because of the way people produce these sounds as they speak, people’s natural vocal apparatus would be unable to pronounce anything between ‘b’ and ‘p’. So when someone hears a sound from the synthetic morphing device, that person tries to compare that sound with what he or she would have to do with his or her voice device to produce this sound. Since a human tuning device can only produce ‘b’ or ‘p’, all continuous synthetic stimuli will be perceived as ‘b’ or ‘p’, whichever is closest.
@@
The hypothesis was also the basis of what is now called the ‘motor theory of speech perception’, which assumes that people perceive spoken words by recognizing the gestures in the speech channel used to pronounce them, rather than by identifying the sound patterns that produce the speech (Liberman et al., 1967). The theory came under fire when it was found that ‘identification’ and ‘discrimination’ of stimuli not at all associated with speech behave in a similar way to stimuli associated with speech when measured in a similar manner (Lane, 1965). It was also found that children, even before they could speak, exhibited the specific categorical perception effect associated with speech perception that had been identified in adults (Eimas et al., 1971).
@@
So, step by step it became clear that categorical perception was a much more general phenomenon than just associated with speech, and the connection was made with earlier findings having to do with the way stimuli are organized. In particular, Lawrence’s experiments and his hypothesis of ‘acquired distinctiveness’ revealed a phenomenon that turned out to be a very basic effect of perception. The hypothesis of acquired distinctiveness states that stimuli for which one is taught to give a different response to them become more distinctive, while stimuli for which one is taught to give the same response to them become more similar (Lawrence, 1949). Both effects are at work in humans in a multitude of perceptions, stimuli that fall within the same category are perceived as more similar, while stimuli that fall into different categories are perceived as more different. What happens with ‘colors’ is a good example of the phenomenon of categorical perception. We see a discrete set of colors while from physical reality there is a continuum of different frequencies presented to us as stimuli. The warping effect of the categorical perception of colors consists in two stimuli that both fall within the category of, for example, green to be perceived more equally than two stimuli one of which falls within the category of green and the second within the category of blue, even if from a physical perspective both pairs of stimuli have the same difference in frequency. The cooperation of these two effects, a contraction within an existing category and a dilation between different categories, causes a clumping of colors, ultimately leading to the colors that we distinguish.
@@
Given the research on colors that we present in this article, it is appropriate to mention Eleanor Rosch’s work on colors. Indeed, it was that work that inspired her to propose the prototype theory of concepts, which remains one of the most important theories of concepts today (Rosch, 1973). The basic idea of prototype theory is that there exists a central element for a concept, which we call the prototype, relative to which the exemplars of the concept can be placed within a graded structure. Rosch suggested the idea that would develop into the primary model for concepts by studying the categorical structure among the Dani for colors and basic shapes. The Dani are a people living in Papua New Guinea, with the peculiarity that they have only two words to denote colors, one meaning Bright and the other Dark  [^1] As in other articles of our Brussels research group, we denote concepts, when they appear as subjects in the text, by writing them beginning with a capital letter and in italic.. The Dani also have no words in their language for basic shapes such as Circle, Square and Triangle. Rosch examined whether there was a difference in learning between two groups of Dani volunteers, with one group learning colors and basic shapes, starting with stimuli that are prototype colors and prototype basic shapes, while the other group learned to start with stimuli that are different distortions of these prototypes. In a significant way, it was found that for both colors and basic shapes, learning was more qualitative for the group taught starting from the prototype stimuli. This evaluation of ‘more qualitative learning’ took into account the three features of how this can be measured, namely by the ease of learning sets of shape categories when a particular type was the prototype, by the ease of learning individual types within sets, and by rank order of rating types as the best example of categories, when the prototypes of both colors and shapes were the stimuli in the learning process.
@@
Mathematical prototype models based on fuzzy sets were developed for concepts and experimentally tested, and it seemed that the way in which by warping, i.e., contraction when stimuli fall into the same category and dilation when they fall into different categories, concepts emerge and grow from stimuli, had finally found a form to understand what is taking place at a fundamental theoretical level (Collier at al., 1973; Rosch, 1975; Rosch et al., 1976; Smith & Medin, 1981; Medin at al., 1984; Geeraerts et al., 2001; Johanden & Kruschke, 2005).
@@
A fundamentally not understood problem however, even when good mathematical models existed by which a concept and its set of exemplars and features could be modeled according to the approach of prototype theory, was the description of the combination of two concepts. This problem was noted in a first publication in which the combination of the concept Pet with the concept Fish served as an example, and therefore the problem of combining two concepts is often called the ‘pet-fish problem’, or also the ‘guppy effect’, because Guppy was the exemplar used to illustrate what goes wrong with prototype theory when concepts are combined (Osherson & Smith, 1981). The question posed was ‘how is it that Guppy is not a typical example of a Pet, nor a typical example of a Fish, but a very typical example of a Pet-Fish’. It is starting from this ‘guppy effect’ that in our Brussels research group it was tried successfully whether a quantum formalism could be used to model the guppy effect mathematically, where then the sudden relevance of Guppy as a typical exemplar would be explained as an ‘interference effect’ between the concepts Pet and Fish when combined (Aerts & Gabora, 2005a, b). These first models using the quantum formalism were further developed and refined in the following years (Aerts, 2009a, b) and their relevance to quantum information science was demonstrated by also modeling data obtained from the World-Wide Web in a similar way with a quantum model (Aerts et al., 2012).
@@
Meanwhile, it became clear that there were many more similarities between the structure and dynamics to which concepts of human language are subject and the structure and dynamics of quantum entities as described by quantum mechanics. The well-studied phenomenon of quantum entanglement, for example, also occurs with concepts, it is possible to combine concepts in a very similar way as is the case with quantum entities, such that experiments on these combinations violate Bell’s inequalities, which is the experimental test for the presence of entanglement (Aerts & Sozzo, 2011, 2014). Another more recent finding consists in identifying the statistical properties of texts of stories, short stories and stories of the length of novels. It could be shown in a very convincing way that the statistics inherent in such texts is of the Bose-Einstein type, hence the same as the statistics of a class of quantum entities called bosons, such as photons. It could also be shown that it cannot be of the Maxwell-Boltzmann type – as one would expect it to be, since Maxwell-Boltzmann is the common statistics for a collection of classical entities (Aerts & Beltran, 2020). It was also investigated and shown that regarding the thermodynamic properties of texts representing stories it is the quantum mechanical von Neumann entropy that is relevant and not the classical Shannon entropy (Aerts & Beltran, 2022a, b). But the most important finding of a recent nature as far as the present article is concerned, is the one related to the phenomenon of quantization. It became clear that the phenomenon of categorical perception triggers a dynamics that leads to the emergence of quanta (Aerts & Aerts Arguëlles, 2022). In the present article, we will again focus on categorical perception and how it gives rise to the formation of quanta, and more specifically, we will show that the mechanism of categorical perception is intrinsically present in the structure of the quantum measurement itself. We will then investigate in what way colors can be considered to be quanta of light for human vision.
@@
Rosch, during her work on colors, assumed that the way humans see colors is determined, probably genetically, at birth, also to regard the color frequencies of the basic colors. More so, that the situation with colors is thus, was actually her inspiration for proposing the prototype theory for concepts. This also means that the so-called Sapir-Whorf hypothesis, namely that there is an influence of ‘how we name categories within a language’ and ‘the perception of stimuli belonging to these categories’, was believed to be not applicable to colors. And this was generally accepted, colors were thought to be not subject to how they were named in different languages. Not only do most cultures divide the groups of colors similarly and give them separate names, but for the few cultures where this is not the case, the areas of compression and dilation were assumed to be the same. It was believed that we all see blue and green in the same way, with a blurred area in between, even if the naming is not the same (Berlin & Kay, 1969). However, this view was challenged by studies that nevertheless identified effects of the words designating colors. Comparative research on colors between speakers of Setswana, a Bantu language spoken by about 8.2 million people in South Africa, and speakers of English found many similarities, but also identified differences relevant in terms of the Sapir-Whorf hypothesis (Davies et al., 1998). Speakers of Berinmo, an indigenous language in Papua New Guinea, have only one word, ‘nol’, for what English speakers call green and blue. The difference they made compared to speakers of English in color discrimination tasks regarding shades between green and blue was investigated and determined (Davidoff, 2001). Later evidence was also found that linguistic categories influence categorical perception primarily in the right visual field. Since the right visual field is controlled by the left hemisphere, this finding was explained by the fact that language skills are also located in the left hemisphere (Regier & Kay, 2009).
@@
More recent experiments, meanwhile, have demonstrated very thoroughly that language and the names given have an influence on the categorization that takes place over very primitive visual perceptions. Nine-month-old infants were shown a continua of new creature-like objects. There was a learning phase in which the infants were shown that objects from one end of the perceptual continuum moved to the left and objects from the other end moved to the right. For one group of infants, the objects were always called by the same name, while for the other group of infants, two different names were used to call the objects depending on whether they belonged to one end or the other of the continuum. The test involved showing new objects from the same continuum to all infants and then seeing if there was a difference between the two groups. What was found is that infants in the one-name condition formed one overarching category and looked at new test objects in either place. Infants in the two-name condition distinguished two categories and correctly anticipated the likely location of test objects even when they were near the poles or near the middle of the continuum (Havy & Waxman, 2016).
@@
However, the experiment with the nine-month-old children we describe above, that demonstrates the effective existence of a Sapir-Whorf effect with colors, is not in contradiction with the findings that inspired Rosch. The colors we customarily distinguish do exist at birth, but the way these colors will play a role dynamically in an individual person’s life, for this more advanced function of what colors are, that is where a Sapir-Whorf effect does play a specific role. As we shall see further in the course of this article, this is the situation that is captured if colors are considered to be quanta of light for human visual perception, indeed, it is then the quantum mechanical structures of superposition and entanglement that can account for these effects.
@@
In her experiments with the Dani, Rosch used the basic colors as defined by Brent Berlin and Paul Kay in their authoritative work ‘Basic Color Terms: Their Universality and Evolution’ (Berlin & Kay, 1969), and we will use these colors as basic colors in this article as well. By the way, let us note that these are not the colors of the rainbow, Berlin & Kay (1969) suggested the following eleven colors as basic colors within the culture where English is used as a native language, White, Black, Red, Yellow, Green, Blue, Brown, Purple, Pink, Orange and Gray. Berlin and Kay’s work studies basic color naming in different language regions and concludes that seven stages in naming these basic colors can be distinguished from an evolutionary perspective. In the first stage, only two colors are named, White (then also meaning Light) and Black (then also meaning Dark). In a second stage, where three names for basic colors exist, Red is systematically added. To the next two stages, Yellow and Green are introduced, sometimes Yellow first and sometimes Green first. At the next stage, Blue is added, and Brown joins the names of basic colors at the stage after that. We have then reached the seventh stage where eight to eleven names exist for basic colors, the names still added being Purple, Pink, Orange and Gray.

### v2

@@ Abstract
We show that colors are light quanta for human visual perception in a similar way as photons are light quanta for physical measurements of light waves. Our result relies on the identification in the quantum measurement process itself of the warping mechanism which is characteristic of human perception. This warping mechanism makes stimuli classified into the same category perceived as more similar, while stimuli classified into different categories are perceived as more different. In the quantum measurement process, the warping takes place between the pure states, which play the role played for human perception by the stimuli, and the density states after decoherence, which play the role played for human perception by the percepts. We use the natural metric for pure states, namely the normalized Fubini Study metric, to measure distances between pure states, and the natural metric for density states, namely the normalized trace-class metric, to measure distances between density states. We then show that when pure states lie within a well-defined region surrounding an eigenstate, the quantum measurement, namely the process of decoherence, contracts the distance between these pure states, while the reverse happens for pure states lying in a well-defined region between two eigenstates, for which the quantum measurement causes a dilation. We elaborate as an example the situation of a two-dimensional quantum measurement described by the Bloch model and apply it to the situation of two colors ‘Light’ and ‘Dark’. We argue that this analogy of warping, on the one hand in human perception and on the other hand in the quantum measurement process, makes colors to be quanta of light for human vision.
Keywords: human vision, categorical perception, quantum measurement, warping, Bloch sphere, quantization, basic colors, qubits, quantum cognition
@@ 1 Introduction
We will argue that colors are quanta of light for human vision in a similar way that photons are quanta of light for a physical measuring device. To gather evidence for this argument, we will draw on various results obtained over the years within the research field of quantum cognition (Aerts & Aerts, 1995; Aerts & Gabora, 2005b; Busemeyer et al., 2006; Aerts, 2009a, b; Bruza & Gabora, 2009; Aerts & Sozzo, 2011; Aerts et al., 2012; Busemeyer & Bruza, 2012; Haven & Khrennikov, 2013; Khrennikov, 2014; Dalla Chiara et al., 2015; Pothos et al., 2015; Blutner & beim Graben, 2016; Moreira & Wichert, 2016; Yearsley, 2017; Aerts Arguëlles, 2018; Busemeyer et al., 2019; Surov et al., 2019; Aerts Arguëlles & Sozzo, 2020; Aerts & Beltran, 2022a), some of which also involve our own work (Aerts Arguëlles, 2018; Aerts Arguëlles & Sozzo, 2020). Nevertheless, we want to make this article understandable to those who have not studied the investigations on which it relies, so that, in addition to references to this research and more technical portions of the article, we will also represent in an intuitive clear way the core aspects in which we frame our claim.
@@
Our argument rests on a fundamental property of human perception, which is not limited specifically to human vision, but is present in all forms of human perception, and which has its origin in a phenomenon called ‘categorical perception’ (Harnad, 1987; Goldstone & Hendrickson, 2010; Aerts & Aerts Arguëlles, 2022). Categorical perception ‘introduces a warping’ of the ‘stimuli’ to the ‘percepts’ in a very specific way. That is, stimuli belonging to the same category are perceived in a way more similar than that shows by a physical measurement of the stimuli. Additionally, for stimuli belonging to different categories, the reverse warping takes place, and they are perceived in a way more different than a physical measurement of the stimuli shows. There is a long history (Liberman et al., 1957, 1967; Lane, 1965; Eimas et al., 1971; Lawrence, 1949; Berlin & Kay, 1969; Davies et al., 1998; Davidoff, 2001; Regier & Kay, 2009; Havy & Waxman, 2016; Hess et al., 2009; Disa et al., 2011; Sidman, 1994; Schusterman et al., 2000; Rosch, 1973; Collier at al., 1973) in which the phenomenon of categorical perception was identified incrementally, and in Section 2 we give a brief description of this genesis.
@@
In the reasoning we develop, we also rely on a detailed study of the quantum measurement process. We have shown in Aerts et al., (2025) that the quantum measurement process involves a warping that is the same as the warping that occurs in categorical perception. We will then gather the arguments that allow us to consider colors as quanta of light for human vision in a similar way that photons are considered quanta of light for physical measurements. For simplicity, we introduce the mathematical elements for our analysis for the case of two colors that we call Light and Dark, mentioning that fundamentally the same analysis can be made for multiple colors with mathematical elements that then belong to a more than two-dimensional complex Hilbert space. No other specific difficulties appear in this more dimensional situation that are not encountered in the two-dimensional case, namely that of two colors and a two-dimensional Hilbert space, which we consider in detail in the present article.
@@
In Section 2 we give a brief overview of the phenomenon of categorical perception with specific attention to the case of colors. In Section 3 we introduce, in a detailed and self-prescriptive way, so that no prior knowledge is necessary, the quantum measurement model. In Section 3, we give an overview of what was shown in Aerts et al., (2025) and present the way that the quantum measurement changes distances between states illustrating that this way contains exactly the warping associated with categorical perception. In other words, the bias caused by human perception is intrinsically contained in the structure of the quantum measurement process. In Section 4 we gather the elements of the previous sections to show our claim that we can consider colors as quanta of light for human vision.
@@ 2 Categorical Perception, Quantization and Colors
In this section we wish to introduce the phenomenon of categorical perception because the nature of this phenomenon, along with the details of the measurement model of quantum mechanics, which our next Section 3 deals with, plays a fundamental role in our claim that colors are quanta of light for human vision. It is interesting to reflect for a moment on the emergence of what was later called categorical perception because it illustrates well the role that certain goals and hypotheses can play.
@@
Around 1950, research on speech perception came into focus as work was being done on a speech device, the ‘pattern playback machine’, which was intended to make it possible to automatically convert texts into spoken form so that, for example, blind people could read with it. This required analyzing speech very thoroughly, and it is within this setting that Alvin Liberman identified the phenomenon that would become known as categorical perception. More specifically, he noticed that when generating a continuum of evenly distributed consonant-vowel syllables with endpoints reliably identified as ‘b’, ‘d’ and ‘g’, there is a point of rapid decrease in the probability of hearing the sound as ‘b’ to hearing it as ‘d’. Later, there is a rapid switch from ‘d’ to ‘g’ (Liberman et al., 1957). Liberman formulated an original hypothesis by which he wished to explain why people perceive an abrupt change between ‘b’ and ‘p’ in the way speech sounds are heard in contrast to what happens with a synthetic morphing device that produces the sounds with a continuous transition. His hypothesis was that this phenomenon is due to a limitation of the human speech apparatus, which, due to the muscular nature of its construction, would be unable to produce continuous transitions. Because of the way people produce these sounds as they speak, people’s natural vocal apparatus would be unable to pronounce anything between ‘b’ and ‘p’. So when someone hears a sound from the synthetic morphing device, that person tries to compare that sound with what he or she would have to do with his or her voice device to produce this sound. Since a human tuning device can only produce ‘b’ or ‘p’, all continuous synthetic stimuli will be perceived as ‘b’ or ‘p’, whichever is closest.
@@
The hypothesis was also the basis of what is now called the ‘motor theory of speech perception’, which assumes that people perceive spoken words by recognizing the gestures in the speech channel used to pronounce them, rather than by identifying the sound patterns that produce the speech (Liberman et al., 1967). The theory came under fire when it was found that ‘identification’ and ‘discrimination’ of stimuli not at all associated with speech behave in a similar way to stimuli associated with speech when measured in a similar manner (Lane, 1965). Children, even before they could speak, were also found to have the specific categorical perception effect associated with speech perception that had been identified in adults (Eimas et al., 1971).
@@
So, step by step, it became clear that categorical perception was a much more general phenomenon than just associated with speech, and the connection was made with earlier findings having to do with the way stimuli are organized. In particular, Lawrence’s experiments and his hypothesis of ‘acquired distinctiveness’ revealed a phenomenon that turned out to be a very basic effect of perception. The hypothesis of acquired distinctiveness states that stimuli for which one is taught to give a different response become more distinctive, while stimuli for which one is taught to give the same response become more similar (Lawrence, 1949). Both effects are at work in humans in a multitude of perceptions, stimuli that fall within the same category are perceived as more similar, while stimuli that fall into different categories are perceived as more different. What happens with ‘colors’ is a good example of the phenomenon of categorical perception. We see a discrete set of colors while from physical reality there is a continuum of different frequencies presented to us as stimuli. The warping effect of the categorical perception of colors consists in two stimuli that both fall within the category of, for example, green to be perceived more equally than two stimuli, one of which falls within the category of green and the second within the category of blue, even if from a physical perspective both pairs of stimuli have the same difference in frequency. The cooperation of these two effects, a contraction within an existing category and a dilation between different categories, causes a clumping of colors, ultimately leading to the colors that we distinguish.
@@
Given the research on colors that we present in this article, it is appropriate to mention Eleanor Rosch’s work on colors. In fact, it was that work that inspired her to propose the prototype theory of concepts, which remains one of the most important theories of concepts today (Rosch, 1973). The basic idea of prototype theory is that there exists a central element for a concept, which we call the prototype, relative to which the exemplars of the concept can be placed within a graded structure. Rosch suggested the idea that would develop into the primary model for concepts by studying the categorical structure among the Dani for colors and basic shapes. The Dani are people who live in Papua-New Guinea, with the peculiarity that they have only two words to denote colors, one meaning Bright and the other Dark  [^1] As in other articles of our Brussels research group, we denote concepts, when they appear as subjects in the text, by writing them beginning with a capital letter and in italic.. The Dani also do not have words in their language for basic shapes such as Circle, Square and Triangle. Rosch examined whether there was a difference in learning between two groups of Dani volunteers, one group learning colors and basic shapes, starting with stimuli that are prototype colors and prototype basic shapes, while the other group learned to start with stimuli that are different distortions of these prototypes. In a significant way, it was found that for both colors and basic shapes, learning was more qualitative for the group taught starting from the prototype stimuli. This evaluation of ‘more qualitative learning’ took into account the three features of how this can be measured, namely the ease of learning sets of shape categories when a particular type was the prototype, by the ease of learning individual types within sets, and by rank order of rating types as the best example of categories, when the prototypes of both colors and shapes were the stimuli in the learning process.
@@
Mathematical prototype models based on fuzzy sets were developed for concepts and experimentally tested, and it appeared that the way in which by warping, i.e. contraction when stimuli fall into the same category and dilation when they fall into different categories, concepts emerge and grow from stimuli had finally found a form to understand what is taking place at a fundamental theoretical level (Collier at al., 1973; Rosch, 1975; Rosch et al., 1976; Smith & Medin, 1981; Medin at al., 1984; Geeraerts et al., 2001; Johanden & Kruschke, 2005).
@@
However, a fundamentally not understood problem, even when good mathematical models existed by which a concept and its set of exemplars and features could be modeled according to the approach of prototype theory, was the description of the combination of two concepts. This problem was noted in a first publication in which the combination of the concept Pet with the concept Fish served as an example, and therefore the problem of combining two concepts is often called the ‘pet-fish problem’, or also the ‘guppy effect’, because Guppy was the exemplar used to illustrate what goes wrong with prototype theory when concepts are combined (Osherson & Smith, 1981). The question posed was ‘How is it that Guppy is not a typical example of a Pet, nor a typical example of a Fish, but a very typical example of a Pet-Fish’. It is starting from this ‘guppy effect’ that in our Brussels research group it was tried successfully whether a quantum formalism could be used to model the guppy effect mathematically, where then the unexpected relevance of Guppy as a typical exemplar would be explained as an ‘interference effect’ between the concepts Pet and Fish when combined (Aerts & Gabora, 2005a, b). These first models using the quantum formalism were further developed and refined in the following years (Aerts, 2009a, b) and their relevance to quantum information science was demonstrated by also modeling data obtained from the World Wide Web in a way similar to a quantum model (Aerts et al., 2012).
@@
Meanwhile, it became clear that there were many more similarities between the structure and dynamics to which concepts of human language are subject and the structure and dynamics of quantum entities as described by quantum mechanics. The well-studied phenomenon of quantum entanglement, for example, also occurs with concepts, it is possible to combine concepts in a very similar way as is the case with quantum entities, such that experiments on these combinations violate Bell’s inequalities, which is the experimental test for the presence of entanglement (Aerts & Sozzo, 2011, 2014). Another more recent finding consists of identifying the statistical properties of text of stories, short stories, and stories of length of novels. It could be shown in a very convincing way that the statistics inherent in such texts are of the Bose-Einstein type, hence the same as the statistics of a class of quantum entities called bosons, such as photons. It could also be shown that it cannot be of Maxwell-Boltzmann type, as one would expect it to be, since Maxwell-Boltzmann is the common statistics for a collection of classical entities (Aerts & Beltran, 2020). It was also investigated and shown that regarding the thermodynamic properties of texts representing stories, it is the quantum mechanical von Neumann entropy that is relevant and not the classical Shannon entropy (Aerts & Beltran, 2022a, b). But the most important finding of a recent nature as far as the present article is concerned is that related to the phenomenon of quantization. It became clear that the phenomenon of categorical perception triggers a dynamics that leads to the emergence of quanta (Aerts & Aerts Arguëlles, 2022). In the present article, we will again focus on categorical perception and how it gives rise to the formation of quanta, and we will make use of the result obtained in Aerts et al., (2025), namely that the mechanism of categorical perception is intrinsically present in the structure of the quantum measurement itself. We will then investigate in what way colors can be considered to be quanta of light for human vision.
@@
Rosch, during her work on colors, assumed that the way humans see colors is determined, probably genetically, at birth, also the color frequencies of the basic colors. More so, the situation with colors being as such was actually her inspiration for proposing the prototype theory for concepts. This also means that the so-called Sapir-Whorf hypothesis, namely that there is an influence of ‘how we name categories within a language’ and ‘the perception of stimuli belonging to these categories’, was believed to be not applicable to colors. This was generally accepted in the aftermath of Rosch’s work on colors. Colors were thought to not depend on what they were called in different languages. Not only do most cultures divide the groups of colors similarly and give them separate names, but for the few cultures where this is not the case, the areas of compression and dilation were assumed to be the same. It was believed that we all see blue and green in the same way, with a blurred area in between, even if the naming is not the same (Berlin & Kay, 1969). However, this view was challenged by studies that nevertheless identified effects of the words designating colors. Comparative research on colors between Setswana speakers, a Bantu language spoken by about 8.2 million people in South Africa, and English speakers found many similarities, but also identified relevant differences in terms of the Sapir-Whorf hypothesis (Davies et al., 1998). The speskers of Berinmo, an indigenous language in Papua New Guinea, have only one word, ‘nol’, for what English speakers call green and blue. The difference they made compared to English speakers in color discrimination tasks with respect to the shades between green and blue was investigated and determined (Davidoff, 2001). Later, evidence was also found that linguistic categories influence categorical perception primarily in the right visual field. Since the right visual field is controlled by the left hemisphere, this finding was explained by the fact that language skills are also located in the left hemisphere (Regier & Kay, 2009).
@@
More recent experiments, meanwhile, have demonstrated very thoroughly that language and the names given have an influence on the categorization that takes place over very primitive visual perceptions. Nine-month-old infants were shown a continua of new creature-like objects. There was a learning phase in which the infants were shown that objects from one end of the perceptual continuum moved to the left and objects from the other end moved to the right. For one group of infants, the objects were always called by the same name, while for the other group of infants, two different names were used to call the objects depending on whether they belonged to one end or the other of the continuum. The test involved showing new objects from the same continuum to all infants and then seeing if there was a difference between the two groups. What was found is that infants in the one-name condition formed one overarching category and looked at new test objects in either place. Infants in the two-name condition distinguished two categories and correctly predicted the likely location of test objects even when they were near the poles or near the middle of the continuum (Havy & Waxman, 2016).
@@
However, the experiment with the nine-month-old children we describe above, which demonstrates the effective existence of a Sapir-Whorf effect with colors, is not in contradiction with the findings that inspired Rosch. The colors we customarily distinguish do exist at birth, but the way these colors will play a role dynamically in an individual person’s life, for this more advanced function of what colors are, is where a Sapir-Whorf effect does play a specific role. As we shall see further in the course of this article, this is the situation that is captured if colors are considered to be quanta of light for human visual perception, indeed, it is then the quantum mechanical structures of superposition and entanglement that can account for these effects.
@@
In her experiments with the Dani, Rosch used the basic colors defined by Brent Berlin and Paul Kay in their authoritative work ‘Basic Color Terms: Their Universality and Evolution’ (Berlin & Kay, 1969), and we will use these colors as basic colors in this article as well. By the way, let us note that these are not the colors of the rainbow, Berlin & Kay (1969) suggested the following eleven colors as basic colors within the culture where English is used as a native language, White, Black, Red, Yellow, Green, Blue, Brown, Purple, Pink, Orange and Gray. Berlin and Kay’s work studies basic color naming in different language regions and concludes that seven stages in naming these basic colors can be distinguished from an evolutionary perspective. In the first stage, only two colors are named, White (then also meaning Light) and Black (then also meaning Dark). In a second stage, where there are three names for basic colors, Red is systematically added. In the next two stages, Yellow and Green are introduced, sometimes Yellow first and sometimes Green first. In the next stage, Blue is added, and Brown joins the names of the basic colors in the next stage. We have then reached the seventh stage where there are eight to eleven names for basic colors, the names still added being Purple, Pink, Orange and Gray.